OpenStudy (anonymous):

find the definite integral ∫(sin^3)6x(cos^4)6xdx

OpenStudy (anonymous):

No domain, indefinite...

OpenStudy (anonymous):

Go to www.WolframAlpha.com Paste in the following to the right of the equal sign: Integrate[Sin[x]^3*6*x*Cos[x]^4*(6*x), x] When the answer is presented, click on "show steps" The integration procedure is very lengthy for this problem. A solution copy from my own Mathematica 8 Home Edition is attached.

OpenStudy (anonymous):

yeah, I'm a little lost by looking at it.

OpenStudy (anonymous):

Maybe u could get away with rewriting the cos^4 as (1-sins^2)^2?

myininaya (myininaya):

\[\int\limits_{}^{}\sin^2(6x)*\cos^4(6x)*\sin(6x)dx\] \[\int\limits_{}^{}(1-\cos^2(6x))\cos^4(6x)*\sin(6x)dx\] \[\int\limits_{}^{}(\cos^4(6x)*\sin(6x)-\cos^2(6x)\cos^4(6x)*\sin(6x))dx\] \[\int\limits_{}^{}\cos^4(6x)*\sin(6x) dx-\int\limits_{}^{}\cos^6(6x)*\sin(6x)dx\] ------------- Let \[u=\cos(6x)=> \frac{du}{dx}=-6\sin(6x) => du=-6\sin(6x) dx\] ------- \[\frac{-1}{6}\int\limits_{}^{}u^4du-\frac{-1}{6}\int\limits_{}^{}u^6du\]

myininaya (myininaya):

\[\frac{-1}{6}*\frac{u^5}{5}+\frac{1}{6}*\frac{u^7}{7}+C\] \[\frac{-1}{6}*\frac{\cos^5(6x)}{5}+\frac{1}{6}*\frac{\cos^7(6x)}{7}+C\]

myininaya (myininaya):

\[-\frac{\cos^5(6x)}{30}+\frac{\cos^7(6x)}{42}+C\]

OpenStudy (anonymous):

Where's the sin 6x come from?

myininaya (myininaya):

sin^3(6x)=sin^2(6x)sin(6x)

OpenStudy (anonymous):

Ah!

OpenStudy (anonymous):

smart

OpenStudy (anonymous):

lol, I'm going to start paying you, myininaya

myininaya (myininaya):

CHECKING: \[[\frac{-\cos^5(6x)}{30}+\frac{\cos^7(6x)}{42}+C]'\] \[\frac{-1}{30}*5*\cos^4(6x)*6*(-\sin(6x))+\frac{1}{42}*7*\cos^6(6x)*6*(-\sin(6x))+0\] \[\cos^4(6x)\sin(6x)-\cos^6(6x)\sin(6x)=\sin(6x)(\cos^4(6x)-\cos^6(6x))\] \[=\cos^4(6x)\sin(6x)(1-\cos^2(6x))=\cos^4(6x)\sin(6x)\sin^2(6x)\] \[=\cos^4(6x)\sin^3(6x)\]

OpenStudy (anonymous):

Isn't it easier to do what I suggested and just integrate Sin^7(6x) -2Sin^5(6x) +Sin^7(6x) ?

myininaya (myininaya):

lol

OpenStudy (anonymous):

Sorry, should be Sin^7(6x) -2Sin^5(6x) +Sin^3(6x)

myininaya (myininaya):

i always try to play with the odd power if there is one

myininaya (myininaya):

how did you get that estudier?

OpenStudy (anonymous):

What I said in my first post use Cos^2 + Sin^2 = 1 to replace Cos^4

OpenStudy (anonymous):

And multiply it out, of course....

OpenStudy (anonymous):

could either of you extraordinary gentlemen perhaps help me solve \[\int\limits_{}^{} \sec^57xtan7xdx\]

myininaya (myininaya):

it might be okay but I think you will still need to us identity to integrate you know so we get du part somewhere

OpenStudy (anonymous):

That's a new question....put it up.

OpenStudy (anonymous):

There's a Tans Secs identity...

OpenStudy (anonymous):

1 + Tan^2 = Sec^2

myininaya (myininaya):

Remember: \[\sin^2x+\cos^2x=1\] this next item can be obtain by dividing the first item by cos^2x \[\tan^2x+1=\sec^2x\] \[\int\limits_{}^{}\sec^2(7x)\tan(7x)\sec^2(7x)dx\] and also remember that (tanx)'=sec^2x \[\int\limits_{}^{}(\tan^2(7x)+1)\tan(7x)\sec^2(7x)dx\] \[\int\limits_{}^{}\tan^2(7x)\tan(7x)\sec^2(7x)dx+\int\limits_{}^{}\tan(7x)\sec^2(7x)dx\] \[\int\limits_{}^{}\tan^3(6x)\sec^2dx+\int\limits_{}^{}\tan(7x)\sec^2(7x)dx\]

myininaya (myininaya):

oops we had sec^5(7x) not sec^4(7x)

OpenStudy (anonymous):

Details...

myininaya (myininaya):

\[\int\limits_{}^{}\sec^4(7x)\sec(7x)\tan(7x)dx\] now what is the derivative of secx? this one was way easier than the first

myininaya (myininaya):

\[u=\sec(7x)=> du=7\sec(7x)\tan(7x)dx\]

myininaya (myininaya):

\[\frac{1}{7}\int\limits_{}^{}u^4 du\]

myininaya (myininaya):

\[\frac{1}{7}*\frac{u^5}{5}+C=\frac{1}{35}u^5+C=\frac{1}{35}\sec^5(7x)+C\]

OpenStudy (anonymous):

why not make u=7x? same thing, but wouldn't it be faster?

OpenStudy (anonymous):

but again, you have my eternal gratitude

myininaya (myininaya):

if you let u=7x, you would have to make another substitution

OpenStudy (anonymous):

oh, right.